A beam of light of wavelength 600 nm from distant source for a single slit 1 mm wide and resulting diffraction pattern is observed on a screen 2cm away. The distance between the First dark fringe on either side of the central Maximum is
1
Expert's answer
2020-09-03T14:14:35-0400

For a dark fringe to be formed, "\frac {dy}{D}=\lambda"

"\lambda=600nm= 600\times 10^{-9}"


"d= 1mm= 10^{-3}"


"D=2m" D=2m


"y=\frac {Dy}{d}= \frac {2\times 600\times10^{-9})}{10^{-3}}= 1.2mm"


The distance between the first dark fringe on either side of the central beam is


"2\times y= 2\times1.2= 2.4mm."


Thus Answer is "2.4mm"